OpenStudy (anonymous):

Which of the following points lies on the circle whose center is at the origin and whose radius is 10?

OpenStudy (whpalmer4):

formula for a circle with center \((h,k)\) and radius \(r\) is \[(x-h)^2+(y-k)^2 = r^2\]As your circle has center at the origin and radius 10, it becomes \[x^2+y^2=10^2\] Try out your points in the formula and see if they work.